OpenStudy (anonymous):

Please helppppppp triple sad face!!!! Elimination method. 5x + 2y = 55 2x - y = 19 answer is x+y=12 ...... I need to know why

OpenStudy (anonymous):

Let Me Guess, You Are HomeschooleD?

OpenStudy (anonymous):

So Aisha, The Elimination method is all about trying to eliminate a variable in one of the equations so we can find the other. So

OpenStudy (anonymous):

i kno how to do the elimination method but this is a trick question

OpenStudy (anonymous):

by observing the equations we can see that by multiplying the 2nd equation by 2 we get 2(2x-y)=2*19; 4x-2y=38, now note the relation between the third equation that we got and the first one. In the first one we have +2y and in the third we have -2y

OpenStudy (anonymous):

but why is the answer x+y=12

OpenStudy (anonymous):

now add the third equation to the first we get 5x+2y\[5x+2y+(4x-2y)=55+(38)\] thus 9x=93, so x=93/9, now we plug in x=93/9 in equation 2 we get y=186/9-19=15/9 now x+y=108/9=12

OpenStudy (anonymous):

wheres the y=186/9-19=15/9 come from?

OpenStudy (anonymous):

@Xesevoli

OpenStudy (anonymous):

@yomamabf

OpenStudy (anonymous):

we got x=93/9 right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

plug it in 2x-y=19

OpenStudy (anonymous):

i got something really weird i dont know where u came across 12

OpenStudy (anonymous):

you agree that x=93/9?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok \[2*(93/9)-y=19\] \[186/9-y=(19*9)/9=171/9\]

OpenStudy (anonymous):

whered the 186 come from?

OpenStudy (anonymous):

\[2*x=2*93/9=(2*93)/9\]

OpenStudy (anonymous):

wait what? whered that 2 come from??????

OpenStudy (anonymous):

wouldnt it be 4(93/9)

OpenStudy (anonymous):

we plugged in x=93/9 in 2x - y = 19

OpenStudy (anonymous):

okay gotcha

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

and then what

OpenStudy (anonymous):

by solving for y, you get y=15/9

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

x+y=93/9+15/9=108/9=(12*9)/9=12

OpenStudy (anonymous):

isnt it -15/9?

OpenStudy (anonymous):

-y=-15/9, so y=15/9

OpenStudy (anonymous):

oh okay gotcha

OpenStudy (anonymous):

sorry im slow so where does the 108 come from

OpenStudy (anonymous):

x+y=93/9+15/9=(93+15)/9=108/9=(12*9)/9=12

OpenStudy (anonymous):

why do u multiply 12 times 9?

OpenStudy (anonymous):

108=12*9 so when we divide by 9 only 12 is left

OpenStudy (anonymous):

im still lost as to where you got that 108 from

OpenStudy (anonymous):

ok i figured it out but where di the 12 come from